Fermi National Laboratory

Accelerator Update for November 10, 2008 - November 12, 2008


Monday Nov 10

The day shift began with the Tevatron (TeV) turn on underway for studies, with Operations monitoring the Recycler stash (169.9E10), with the Antiproton Source stacking (83.9E10), with the Pelletron in access, and with the NuMI, MiniBooNE, and Meson MTest experiments taking beam.

At 9:16 AM, Operations reported that everything was ready to turn on the TeV except for the sector A3 VCB. An expert reported that its breaker had to be replaced.

A RF expert began working on station MIRF17 at 11:24 AM.

At 11:47 AM, the TeV quenched at sector A1 during its first ramp. Experts investigated.

A MI expert began a study at 2:15 PM.

At 2:40 PM, the TeV quenched again at sector A1 during its first ramp. Experts investigated. Operators reported that the vacuum at A1 was good.

Preaccelerator experts made a 10-minute access at 2:52 PM into the H- Source to determine the cause of some sparking.

Operations reported at 3:10 PM that they had begun preparing the TeV sector A15 for access.

A MI expert reported at 5 PM that she had completed her MI study.

Recycler experts reported at 6:13 PM that they had completed their work and begun refilling the Pelletron with its spark retardant gas (SF6).

The TeV access into sector A1 was completed at 7:32 PM.

Operators ramped up the TeV successfully and then put it through a dry squeeze at 10:14 PM.

Tuesday Nov 11

The midnight shift began with the TeV in a dry squeeze, with Operations monitoring the stash (163.9E10), with the Pbar stacking (180.8E10), and with NuMI and MiniBooNE taking beam.

Operators reported at 1 AM that they had search and secured the Pelletron (MI-31).

A Recycler expert reported at 2:33 AM that the Pelletron’s Electron Cooling (ECOOL) system was operating normally.

Operations put the TeV through a dry squeeze at 3:01 AM.

Operations established beam to the MTest experiment at 5:49 AM.

Operators reported at 6:51 AM that they had successfully injected protons into the TeV.

At 7:22 AM, Operations reported that a Pbar Debuncher septum (D:ISEPV) had tripped off several times on a ground fault indication. It reset.

At 7:42 AM, Operations reported that the TeV sector C4 BPM Front End (FE) had failed. Operators could not reset the FE. They called in a controls expert.

Operations reported at 8:03 AM that the Recycler had stashed antiprotons twice during the midnight shift.

Linac experts reported at 10:07 AM that they were beginning a study.

At 12:08 PM, the D:ISEPV tripped off on a ground fault indication. It reset.

Operations put the TeV into shot setup at 1:40 PM.

Operations put Pbar into standby mode at 1:50 PM.

Linac experts ended their study around 4 PM.

Operations established store 6553 at 4:51 PM with an initial luminosity of 235.95E30.

At 4:54 PM, Operations reported a problem with the TeV sector E17 Schottky system. Experts investigated.

At 5:47 PM, Operations reported that the Meson Critical Device Controller tripped off for no obvious reason. Operators investigated.

Operations disabled the MTest beam at 6:04 PM.

Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.

Wednesday Nov 12

The midnight shift began with Operations monitoring store 6553 (90.18E30) and the stash (277.7E10), with Pbar stacking (18.7E10), with the Collider Detector Facility (CDF) and D-Zero (D0) using the colliding beams, and with NuMI and MiniBooNE taking beam.

Operations terminated store 6553 at 4:29 AM.

Operations put the TeV into shot setup at 4:31 AM.

Operations established store 6554 at 6:26 AM with an initial luminosity of 300.9E30.

Water system experts reported at 7:04 AM that they had added water to the MI LCW system.

Operations reported at 7:57 AM that the Recycler had stashed six times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac, Booster, and MI experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/12/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory